OpenStudy (anonymous):

A triangle drawn on a map has sides of lengths 7.0 cm, 11 cm, and 13 cm. The shortest of the corresponding real-life distances is 91 km. Find the longest of the real-life distances. Round your answer to the nearest whole number.

OpenStudy (anonymous):

"The shortest of the corresponding real-life distances is 91 km" means that the shortest out of the three measurements (the 7cm, 11cm, and 13cm) is equivelient to 91 km. This means that: 7cm = 91 km in real life then how much does 1 cm = in real life? if we can find that out then we can find how much both the 11 and 13 equivelant to. to find how many km is 1cm we do the following: \[91/7 = 13\] hence, 1 cm = 13 km so 11 cm: 11*13= 143 km and 13 cm = 13*13= 169 km therefore the longest of the real-life distances is 169 km.

OpenStudy (anonymous):

thanks :)